OpenStudy (anonymous):

Medal for the best answer ! Jonah needs to purchase 30 juice packs for his class. While shopping, Jonah discovers the following prices for comparable juice packs. Which offers the best unit price? A. 15 Packs for $3.95 B. 5 Packs for $2.20 C. 20 packs for $6.20 D. 10 packs for $3.20 Is it C??

OpenStudy (haichi):

A=7.9 D=9.60 C=9.60 B=13.2

OpenStudy (haichi):

No medal?

OpenStudy (bibby):

it's C. what are those numbers @Haichi

OpenStudy (anonymous):

Yeah I don't understand where they came from...?

OpenStudy (haichi):

Well D is 3.20 for ten packs so 3.20 30=answer

OpenStudy (anonymous):

?

OpenStudy (haichi):

nvm *sigh*

OpenStudy (bibby):

What I did was get b and d in terms of c. b - multiply both by 4 - 20/8.8 d - multiply both by 2 - 20/12.80 they're both worse deals then a = 15/3.95 = 3.797 =per unit then c = 20/6.20= 3.22 per unit

OpenStudy (haichi):

ok shore